OpenStudy (anonymous):

I'm suppose to solve this equation algebraically. Can anyone help me with solve it? log4(x-3)=6+log4(x) @Algebra

jimthompson5910 (jim_thompson5910):

log4(x-3)=6+log4(x) log4(x-3)-log4(x)=6 log4[(x-3)/(x)] = 6 (x-3)/(x) = 4^6 (x-3)/(x) = 4096 x-3 = 4096x -3 = 4096x-x -3 = 4095x -3/4095 = x -1/1365 = x x = -1/1365

OpenStudy (anonymous):

Thanks!